OpenStudy (christos):

Help me solve 0 < 2/x < 3/5 Medal will be given

OpenStudy (ash2326):

\[0<\frac 2 x< \frac 35\] Multiply by x the whole equation. There is a catch here, we need to check if x is positive or negative \we have \[\frac 2 x >0\] so \[x>0\] Now we can multiply the whole equation by x \[0<2<3x\] from this we have \[2<3x\] Do you get this @Christos?

OpenStudy (ash2326):

oops \[2< \frac {3x}5\]

OpenStudy (anonymous):

\[0<\frac{ 2 }{ x }<\frac{ 3 }{ 5 }\] \[3 x>10\] \[x>\frac{ 10 }{ 3 }\]

OpenStudy (ash2326):

@Christos Do you get this?

OpenStudy (christos):

Yes thank you all! some_someone can you explain me how did you get from step 1 to step 2? did you multiply x with 3 and 2 with 5 and removed the 0? can you do that?

OpenStudy (anonymous):

so you have to get x by itself.

OpenStudy (christos):

so this 0 > 2/x > 3/5 can be writen like this? 0 > 3x > 10 ? Can I keep the 0 at the left?